# Deep learning approach for identification of Hii regions during reionization in 21-cm observations – III. image recovery

### v1

@@ 4.3 Statistical Evaluation of SERENEt
In the left panel of Figure 6, we show the $R^{2}$ score (Equation 3) against the true neutral fraction, $\overline{x}_{\mathrm{HI,true}}$. Each point represents a 21-cm image at a given redshift, as the colorbar indicates. As expected, the images at the lowest ($z\sim 7$) and highest ($z\sim 11$) redshifts show a decline in the $R^{2}$ score due to the inefficiency of PCA method modelling the foreground at the boundaries of the lightcone data that we consider. The average accuracy on the entire random dataset is approximately $R^{2}\right>=0.754$. Meanwhile, the blue dash-dotted line indicates the mean score as a function of volume-averaged neutral fraction, $R^{2}(x_{\mathrm{HI}})$. In Table 1, we show the accuracy of the recovered 21-cm maps and the corresponding predicted binary prior to the different reionization milestones; in bold, we highlight the bin corresponding to the visual example of the fiducial model. The contour lines show the $68$ per cent (green solid line) and $95$ per cent (green dashed line) confidence intervals of the data, with a corresponding average accuracy of $R^{2}\right>=0.763$ and $R^{2}\right>=0.760$, respectively.
In the right panel of Figure 6, we show the correlation plot between RecU-Net’s accuracy in recovering the 21-cm signal ($R^{2}$) and its dependency on the quality of the binary prior map predicted by SegU-Net ($r_{\phi}$), which is defined by Equation 4. The colour of the data points indicates the ground truth neutral fraction at intervals of $0.1$. A dashed black line represents the expected behaviour of the $R^{2}-r_{\phi}$ relation, as we expect that the quality of the binary prior will affect the prediction of the 21-cm signal. Ideally, most points would be located at the top right of the plot. In our case, $68$ per cent of the recovered 21-cm images have $r_{\phi}\geq 0.80$ and $0.45\leq R^{2}\leq 0.92$, indicating that most of the 21-cm signals from different stages of the reionization history are accurately recovered. The $95$ per cent contour includes predictions with overall lower performance, with values $r_{\phi}\geq 0.70$ and $R^{2}\leq 0.97$ for RecU-Net’s predictions.
@@ 4.4 Recovered cylindrical Power Spectra
Figure 8: Left panel: Similar analysis for $R^{2}$ versus volume-averaged neutral fraction, but the score is calculated on 2D power spectra for the random dataset. Right panel: comparison of the score computed in the Fourier space, $R^{2}(P_{k_{\perp}k_{\parallel}})$, and the spatial domain $R^{2}(\delta T_{b})$. We randomly select a $100$ redshift within the available range, $117<z<$ 11, with a frequency depth of ±$\pm 20\rm\,MHz$. The colorbar indicates the central redshift, $z_{c}$. The extremes of the redshift range are not sampled due to the required frequency depth for calculating the power spectra.
@@
Figure 9: Similar visualisation as in Figure 4 comparing different training scenarios applied to the EOS data for different reionization milestones mentioned in the panel title in the leftmost column.
@@
The same features at $k_{\parallel}\approx 0.55$, $1.05$ and $1.60\,\rm Mpc^{-1}$ are present in the case of the no prior and realistic prior, whereas they are less evident in the third case. This is because, in the perfect prior, the recovered 21-cm signal appears to be smoothed and presents fewer fluctuations within the 21-cm regions. Moreover, the large scale mode, at small $k_{\parallel}$ and $k_{\perp}$, appears to be the predominant factor influencing the score estimation, $R^{2}$.
@@
An important conclusion is that an almost perfect reconstruction of the power spectra, $R^{2}\geq 0.95$, does not imply a recovery of the 21-cm signal of the same degree of accuracy, as it is the case for the no prior test. Thus, in future observation, we can expect that techniques that aim to mitigate foreground contamination exclusively in the visibility space require additional steps in the image domain to recover the corresponding 21-cm images for astrophysical and cosmological inferences.
@@
We perform a similar analysis, as shown in Figure 6, but we apply the score to the 2D power spectra for the $300$ realization of the random dataset for the realistic prior case. For each realization of the random dataset, we randomly choose $100$ redshifts as central redshift, $z_{c}$, and we then select a ±$\pm 10\rm MHz$ region around this redshift, from which we calculate $P(k_{\perp},\,k_{\parallel})$. The results in the left panel of Figure 8 confirm that a high score with $R^{2}>0.90$ is achieved by $68$ per cent of the data, while the $95$ per cent contours show results within $R^{2}>0.60$. Similarly to the spatial domain case shown in §4.3, the accuracy decreases for samples close to the redshift limits. Overall, we can expect SERENEt to perform on the same level of accuracy as the average score for the entire random dataset, which is $R^{2}\right>=0.934$. In this analysis, we have fewer or no data close to the redshift limit, which limits to data with reionization history with $\overline{x}_{\rm HI,\,true}\leq 0.95$. This is due to the fact that the power spectra are calculated for a sub-sample of the entire lightcone, centred at a given redshift and with a frequency depth.
@@ 4.5 Recover 21-cm signal on Large Field of View
In Figure 10, we confirm that the two scenarios with prior have predictions with similar and, at the same time, higher accuracy than the no prior results with a difference of $5-10\%$. The redshift evolution of the $R^{2}$ score shows the characteristic depression at the redshift limits of the subvolume due to the pre-processing. The perfect prior shows a wider redshift range of efficiency with lower and upper limits of $z\geq 7.15$ and $z\leq 8.1$. Surprisingly, the no prior has the same upper limit at the cost of accuracy drops at higher redshift, $z\sim 7.35$. Moreover, the accuracy of the realistic prior decreases swifter than the other two cases, so by $z\sim 7.8$, the no prior provides better prediction. The reason could be related to the fact that at this redshift the simulation still presents a few cold, $T_{\mathrm{S}}<T_{\mathrm{\gamma}}$, ionized regions that could confuse the segmentation step, as SegU-Net is trained on data that assumes that only regions with no signal, $\delta T_{b}=0\,\rm mK$, are indeed ionized $x_{\mathrm{HI}}=0$. This misleading does not affect the perfect prior at the contrary of the realistic prior, as it does not require SegU-Net to infer the binary prior from the residual image. On the other hand, the drop in accuracy at higher redshift for the no prior case suggests that, in this case, without the help of the prior, the network mistakes foreground residual as a 21-cm signal.
@@ 4.6 Galaxy Surveys as the Prior Information
We intend to understand if the proposed architecture (illustrated in Figure 2) could work on a prior based on the angular position and the redshift of these ionizing sources. However, this approach requires some understanding of the sources proprieties and their effect on the surrounding IGM. However, in this test, we assume we know the mean-free path of UV photons at high redshift and consider a redshift-dependent relation with two free parameters:
@@
Here, we consider the values of the free parameters $\alpha_{0}=-2.55$ and $z_{0}=2.47$ (Choudhury, 2009). To be able to apply this method, we employed the publicly available version of 21cmFAST_v4555https://github.com/21cmfast/21cmFAST/tree/v4-prep (Davis et al., in prep.). Contrary to its previous publication, this version can generate halo catalogues identified from the Lagrangian initial conditions using an ellipsoidal collapsed barrier (Sheth & Tormen, 2002) and then move them to real space using 2LPT.
@@
To paint an approximate bubble H II bubble around sources, we follow the approach proposed by Korber et al. (2023). Based on the position of the sources and their redshift, we apply Equation 5 and define a spherical region around each source, and we expect the size of the H II bubble to be within this distance from the source. In Figure 12, we show an example of the galaxy prior. The colour map indicates the combined overlap for all the sources within the coeval cube, while red dots are the corresponding halo. The size of the red circles is indicative of the halo mass. In green, we show an example of the spherical region described bu Equation 5 for the most significant source in the simulation. We create a binary prior from the data with overlapping spheres by taking the average value as threshold and labelling pixels with value 00 when the number of overlaps is above the average instead of $1$ if it is below. We refer to this new prior as the galaxy prior. We show the corresponding example of the constructed prior map at redshift $z=9.9$ in Figure 12 with cyan-coloured contours. In this case, the threshold is of $\sim 42$ overlaps.
We employ the resulting galaxy prior maps as the prior input for SERENEt as an alternative case of the realistic prior. Therefore, without retraining the network, we show the resulting prediction of the 21-cm signal in Figure 13 at redshift $z=9.9$, $8.8$ and $7.9$ corresponding to $\overline{x}_{\mathrm{HI,true}}=0.8$, $0.5$ and $0.1$. The last column shows the results from the galaxy prior. On each panel that uses a prior, we over-plot the contour to show the binary prior map employed by the network.
Here, the no prior has a maximum average score of $R^{2}=0.55$ around redshift $z\approx 8$ to then decrease around $R^{2}=0.47$ for $z\approx 9$ and higher redshift. On the other hand, as illustrated in §4.1 and 4.2, the perfect prior provide the best results with an almost constant accuracy of $R^{2}\sim 0.94$ for $z\lesssim 9$ and $R^{2}=0.81$ at higher redshift. It also shows similar features in the predicted 21-cm images, with virtually no signal fluctuations within H I regions. Moreover, once again, the realistic prior provides better results than the no prior case with almost up to twice more accuracy at $z=7.9$, $R^{2}=0.87$, to the same level of accuracy, $R^{2}\sim 0.60$, at higher redshift, $z\gtrsim 9.5$, as discovered in the previous analysis with Figure 5.
[…]
@@ 5 Conclusion
At first, in §4.1, we show an example of the input and output of SERENEt. We then focus on the fiducial model and visually evaluate the recovered 21-cm tomographic data in §4.2. We assess the effectiveness of employing binary priors maps derived from the simulated volume-average neutral fraction maps, $x_{\mathrm{HI}}(\mathbf{r},z)$. Our first conclusion is that the prior has a tangible effect in removing residual foreground with differential brightness of $\delta T_{b}\sim 5\,\rm mK$ that otherwise is mistakenly detected as 21-cm signals. Moreover, the 21-cm signal recovered with a realistic prior visually provides more legitimate results when compared to the perfect prior.
@@
In §4.3, we statistically quantify the prediction on a large dataset that we employ for testing. Our results show that SERENEt gives consistently better reconstructed 21-cm images when the binary prior is included as an additional input. In particular, the realistic prior shows that an average score of $R^{2}\right>=0.75$ on the entire random dataset, with performances that vary between $R^{2}\simeq 0.88$ when $\overline{x}_{\mathrm{HI}}=0.1$ and $R^{2}\simeq 0.73$ when $\overline{x}_{\mathrm{HI}}=0.9$, as shown by the analysis in Table 1. Predictions on the binary maps follow a similar trend with $r_{\phi}\approx 0.73$ at $\bar{x}_{\mathrm{HI}}=0.1$ and $r_{\phi}\approx 0.92$ at $\bar{x}_{\mathrm{HI}}=0.9$. From the recovered 21-cm tomographic dataset, we calculate 2D power spectra $P(k_{\perp},k_{\parallel})$, the analysis in §4.4 shows that the corresponding summary statistic showing an astonishing average score of $R^{2}\right>=0.93$ on the full random dataset.
@@
Moreover, in §4.5, we apply SERENEt to a mock observation with box size of $1.6\,\rm Gpc$, which has a corresponding field of view of approximately $10\,\rm deg$, conform with expected SKA-Low observations. We decompose the large images into smaller ones that fit SERENEt input shape and compared the abovementioned three scenarios. Our network can recover the 21-cm from large FoV SKA-Low mock observations. Moreover, the realistic prior provides better results with higher accuracy than the other cases, indicating that the network trained with the perfect prior relies excessively on the binary prior, as the recovered 21-cm maps have almost no fluctuations in the regions emitting signals. Thus, imperfections in the binary prior give a more balanced training, relying more on the residual image and more stable and accurate predictions. Moreover, we notice that the actual limitation in the redshift range on which we can apply SERENEt is mostly related to limitations in the efficiency of the preprocessing as well as the spin-saturated approximation, $T_{\mathrm{S}}\gg T_{\mathrm{CMB}}$, that we impose on our datasets.
@@
Contrary to our initial intuition, one of our significant discoveries is that training our network on a more realistic prior appears to provide more stable and, in some cases, more accurate predictions. This conclusion was first evident from the results in §4.5, and later it is confirmed in §4.6, where we do a final test and run an additional 21-cm simulation that also provides a galaxy catalogue. From the location of the sources, their redshift and some assumption on the ionizing photons mean-free path, $\lambda_{\mathrm{mfp}}$, we can compute a different binary prior on the assumption that reionization is an inside-out process and thus not assuming that a region void of signal, $\delta T_{b}=0\,\rm mK$ is ionized, i.e. $x_{\mathrm{HI}}(\mathbf{r},z)=0$.
@@
There have been several attempts to address foreground residuals in 21-cm observations (e.g. Gagnon-Hartman et al., 2021; Kennedy et al., 2024). However, these efforts have not focused on reconstructing the full 21-cm signal image data, which contains valuable non-Gaussian information. With SERENEt, we can explore higher-order summary statistics, such as the bispectrum (e.g. Majumdar et al., 2018; Watkinson et al., 2019) and improve our understanding of the underlying astrophysical and cosmological processes that affects the signal (e.g. Jennings et al., 2020; Hutter et al., 2020; Giri & Schneider, 2022; Schaeffer et al., 2023; Schneider et al., 2023).

### v2

@@ 4.3 Statistical Evaluation of SERENEt
In the left panel of Figure 6, we show the $R^{2}$ score (Equation 3) against the true neutral fraction, $\overline{x}_{\mathrm{HI,true}}$. Each point represents a 21-cm image at a given redshift, as the colorbar indicates. As expected, the images at the lowest ($z\sim 7$) and highest ($z\sim 11$) redshifts show a decline in the $R^{2}$ score due to the inefficiency of PCA method modelling the foreground at the boundaries of the lightcone data that we consider. The average accuracy on the entire random dataset is approximately $\left<R^{2}\right>=0.754$. Meanwhile, the blue dash-dotted line indicates the mean score as a function of volume-averaged neutral fraction, $R^{2}(x_{\mathrm{HI}})$. In Table 1, we show the accuracy of the recovered 21-cm maps and the corresponding predicted binary prior to the different reionization milestones; in bold, we highlight the bin corresponding to the visual example of the fiducial model. The contour lines show the $68$ per cent (green solid line) and $95$ per cent (green dashed line) confidence intervals of the data, with a corresponding average accuracy of $\left<R^{2}\right>=0.763$ and $\left<R^{2}\right>=0.760$, respectively.
In the right panel of Figure 6, we show the correlation plot between RecU-Net’s accuracy in recovering the 21-cm signal ($R^{2}$) and its dependency on the quality of the binary prior map predicted by SegU-Net ($r_{\phi}$), which is defined by Equation 4. The colour of the data points indicates the ground truth neutral fraction at intervals of $0.1$. A dashed black line represents the expected behaviour of the $R^{2}-r_{\phi}$ relation, as we expect that the quality of the binary prior will affect the prediction of the 21-cm signal. Ideally, most points would be located at the top right of the plot. In our case, $68$ per cent of the recovered 21-cm images have $r_{\phi}\geq 0.65$ and $0.52\leq R^{2}\leq 0.92$, indicating that most of the 21-cm signals from different stages of the reionization history are accurately recovered. The $95$ per cent contour includes predictions with overall lower performance, with values $r_{\phi}\geq 0.55$ and $R^{2}\leq 0.97$ for RecU-Net’s predictions.
@@ 4.4 Recovered cylindrical Power Spectra
Figure 8: Left panel: Similar analysis for $R^{2}$ versus volume-averaged neutral fraction, but the score is calculated on 2D power spectra for the random dataset. Right panel: comparison of the score computed in the Fourier space, $R^{2}(P_{k_{\perp}k_{\parallel}})$, and the spatial domain $R^{2}(\delta T_{b})$. We randomly select a $100$ redshift within the available range, $7<z<11$, with a frequency depth of $\pm 20\rm\,MHz$. The colorbar indicates the central redshift, $z_{c}$. The extremes of the redshift range are not sampled due to the required frequency depth for calculating the power spectra.
@@
Figure 9: Similar visualisation as in Figure 4 comparing different training scenarios applied to the EOS data for different reionization milestones mentioned in the panel title in the leftmost column. In each plot, we show an inset plot that zooms in on a region of size 256 Mpc per side, corresponding to the dimensions of the network input.
@@
The same features at $k_{\parallel}\approx 0.55$, $1.05$ and $1.60\,\rm Mpc^{-1}$ are present in the case of the no prior and realistic prior, whereas they are less evident in the third case. These features are spectral leakage due to sharp boundaries in the data when computing the 2D power spectrum (e.g., Barry et al., 2019; Wilensky et al., 2021). This sharp discontinuity is stronger along the $k_{\parallel}$ direction due to evolution of the 21-cm signal along the line-of-sight, namely the light-cone effect (e.g., Datta et al., 2012; Giri et al., 2018a). In the perfect prior, the recovered 21-cm signal appears to be smoothed and presents fewer fluctuations within the 21-cm regions. Moreover, the large-scale mode, at small $k_{\parallel}$ and $k_{\perp}$, appears to be the predominant factor influencing the score estimation, $R^{2}$. The amplitude of the signal in the 2D power spectra is sensitive to the global average neutral fraction at the corresponding central redshift, $P_{\mathrm{21cm}}(k_{\perp},k_{\parallel})\propto T^{2}_{0}\,\overline{x}^{2}_{\mathrm{HI}}$ (Lidz et al., 2007; Georgiev et al., 2022). Therefore, in our case, although the $R^{2}$ score is not capturing the spectral leakage effect, it indicates how well the neural network captures the reionization history.
@@
An important conclusion is that an almost perfect reconstruction of the power spectra, $R^{2}\geq 0.95$, does not imply a recovery of the 21-cm signal of the same degree of accuracy, as is the case for the no prior test. Thus, in future observations, we can expect that techniques that aim to mitigate foreground contamination exclusively in the visibility space require additional steps in the image domain to recover the corresponding 21-cm images for astrophysical and cosmological inferences.
@@
We perform a similar analysis, as shown in Figure 6, but we apply the score to the 2D power spectra for the $300$ realization of the random dataset for the realistic prior case. For each realization of the random dataset, we randomly choose $100$ redshifts as central redshift, $z_{c}$, and we then select a $\pm 10\rm MHz$ region around this redshift, from which we calculate $P(k_{\perp},\,k_{\parallel})$. The results in the left panel of Figure 8 confirm that a high score with $R^{2}>0.90$ is achieved by $68$ per cent of the data, while the $95$ per cent contours show results within $R^{2}>0.60$. Similarly to the spatial domain case shown in §4.3, the accuracy decreases for samples close to the redshift limits. Overall, we can expect SERENEt to perform on the same level of accuracy as the average score for the entire random dataset, which is $\left<R^{2}\right>=0.934$. In this analysis, we have fewer or no data close to the redshift limit, which limits the data to reionization history with $\overline{x}_{\ rmHI,\ true}\leq 0.95$. This is due to the fact that the power spectra are calculated for a sub-sample of the entire lightcone, centred at a given redshift and with a frequency depth.
@@ 4.5 Recover 21-cm signal on Large Field of View
In Figure 10, we confirm that the two scenarios with prior have predictions with similar and, at the same time, higher accuracy than the no prior results with a difference of $5-10\%$. The redshift evolution of the $R^{2}$ score shows the characteristic depression at the redshift limits of the subvolume due to the pre-processing. The perfect prior shows a wider redshift range of efficiency with lower and upper limits of $z\geq 7.15$ and $z\leq 8.1$. Surprisingly, the no prior has the same upper limit at the cost of accuracy drops at higher redshift, $z\sim 7.35$. Moreover, the accuracy of the realistic prior decreases more quickly than the other two cases, so by $z\sim 7.8$, the no prior provides better prediction. The reason could be related to the fact that at this redshift the simulation still presents a few cold, $T_{\mathrm{S}}<T_{\mathrm{CMB}}$, ionized regions that could confuse the segmentation step, as SegU-Net is trained on data that assumes that only regions with no signal, $\delta T_{b}=0\,\rm mK$, are indeed ionized $x_{\mathrm{HI}}=0$. This misleading does not affect the perfect prior, in contrast to the realistic prior, as it does not require SegU-Net to infer the binary prior from the residual image. On the other hand, the drop in accuracy at higher redshift for the no prior case suggests that, in this case, without the help of the prior, the network mistakes foreground residual as a 21-cm signal.
@@ 4.6 Galaxy Surveys as the Prior Information
We intend to understand if the proposed architecture (illustrated in Figure 2) could work on a prior based on the angular position and the redshift of these ionizing sources. However, this approach requires some understanding of the source properties and their effect on the surrounding IGM. However, in this test, we assume we know the mean-free path of UV photons at high redshift and consider a redshift-dependent relation with two free parameters:
@@
Here, we consider the values of the free parameters $\alpha_{0}=-2.55$ and $z_{0}=2.47$ (Choudhury, 2009). To be able to apply this method, we employed the publicly available version of 21cmFAST_v4444https://github.com/21cmfast/21cmFAST/tree/v4-prep (Davis et al., in prep.). Contrary to its previous publication, this version can generate halo catalogues identified from the Lagrangian initial conditions using an ellipsoidal collapsed barrier (Sheth & Tormen, 2002) and then move them to real space using 2LPT.
@@
To paint an approximate bubble Hii bubble around sources, we follow the approach proposed by Korber et al. (2023). Based on the position of the sources and their redshift, we apply Equation 5 and define a spherical region around each source, and we expect the size of the Hii bubble to be within this distance from the source. In Figure 12, we show an example of the galaxy prior. The colour map indicates the combined overlap for all sources within the coeval cube, while the red dots correspond to the halo. The size of the red circles is indicative of the halo mass. In green, we show an example of the spherical region described by Equation 5 for the most significant source in the simulation. We create a binary prior from the data with overlapping spheres by taking the average value as a threshold and labelling pixels with value 0 when the number of overlaps is above the average instead of $1$ if it is below. We refer to this new prior as the galaxy prior. We show the corresponding example of the constructed prior map at redshift $z=9.9$ in Figure 12 with cyan-coloured contours. In this case, the threshold is of $\sim 42$ overlaps.
We employ the resulting galaxy prior maps as the prior input for SERENEt as an alternative case of the realistic prior. Therefore, without retraining the network, we show the resulting prediction of the 21-cm signal in Figure 13 at redshift $z=9.9$, $8.8$ and $7.9$ corresponding to $\overline{x}_{\mathrm{HI,true}}=0.8$, $0.5$ and $0.1$. The last column shows the results from the galaxy prior. On each panel that uses a prior, we overplot the contour to show the binary prior map employed by the network.
Here, the no prior has a maximum average score of $R^{2}=0.55$ around redshift $z\approx 8$ to then decrease around $R^{2}=0.47$ for $z\approx 9$ and higher redshift. On the other hand, as illustrated in §4.1 and 4.2, the perfect prior provides the best results with an almost constant accuracy of $R^{2}\sim 0.94$ for $z\lesssim 9$ and $R^{2}=0.81$ at higher redshift. It also shows similar features in the predicted 21-cm images, with virtually no signal fluctuations within Hi regions. Moreover, once again, the realistic prior provides better results than the no prior case with almost up to twice more accuracy at $z=7.9$, $R^{2}=0.87$, to the same level of accuracy, $R^{2}\sim 0.60$, at higher redshift, $z\gtrsim 9.5$, as discovered in the previous analysis with Figure 5.
Finally, the newly introduced prior also provides better results than the simple no prior case even though the prior in this scenario is far from flawless compared to the other two cases, as shown by the contours in Figure 13. The galaxy prior has an accuracy of $R^{2}\sim.75$ compared to the $R^{2}\sim.50$ of the no prior case at $z\approx 8$ and similarly to the previous case, its accuracy decreases at higher redshift to an average of $R^{2}\sim 0.50$. Moreover, as expected, the galaxy prior performs slightly worse than the realistic prior, suggesting that retraining the network on the new prior with a more appropriate $\lambda_{\mathrm{mfp}}$ could vastly improve its performance.
@@ 5 Conclusion
At first, in §4.1, we show an example of the input and output of SERENEt. We then focus on the fiducial model and visually evaluate the recovered 21-cm tomographic data in §4.2. We assess the effectiveness of employing binary prior maps derived from the simulated volume-average neutral fraction maps, $x_{\mathrm{HI}}(\mathbf{r},z)$. Our first conclusion is that the prior has a tangible effect in removing residual foreground with a differential brightness of $\delta T_{b}\sim 5\,\rm mK$, which would otherwise be mistakenly detected as 21-cm signals. Moreover, the 21-cm signal recovered with a realistic prior visually provides more legitimate results when compared to the perfect prior.
@@
In §4.3, we statistically quantify the prediction on a large dataset that we employ for testing. Our results show that SERENEt gives consistently better reconstructed 21-cm images when the binary prior is included as an additional input. In particular, the realistic prior shows that an average score of $\left<R^{2}\right>=0.75$ on the entire random dataset, with performances that vary between $R^{2}\simeq 0.88$ when $\overline{x}_{\mathrm{HI}}=0.1$ and $R^{2}\simeq 0.73$ when $\overline{x}_{\mathrm{HI}}=0.9$, as shown by the analysis in Table 1. Predictions on the binary maps follow a similar trend with $r_{\phi}\approx 0.73$ at $\bar{x}_{\mathrm{HI}}=0.1$ and $r_{\phi}\approx 0.92$ at $\bar{x}_{\mathrm{HI}}=0.9$. From the recovered 21-cm tomographic dataset, we calculate the 2D power spectra $P(k_{\perp},k_{\parallel})$. The analysis in §4.4 shows an astonishing average score of $\left<R^{2}\right>=0.93$ on the random dataset.
@@
Moreover, in §4.5, we apply SERENEt to a mock observation with a box size of $1.6\,\rm Gpc$, which has a corresponding field of view of approximately $10\,\rm deg$, consistent with expected SKA-Low observations. We decompose the large images into smaller ones that fit the SERENEt input shape and compare the abovementioned three scenarios. Our network can recover the 21-cm from large FoV SKA-Low mock observations. Moreover, the realistic prior provides better results with higher accuracy than the other cases, indicating that the network trained with the perfect prior relies excessively on the binary prior, as the recovered 21-cm maps have almost no fluctuations in the regions emitting signals. Thus, imperfections in the binary prior result in a more balanced training, relying more on the residual image and yielding more stable and accurate predictions. Moreover, we notice that the actual limitation in the redshift range on which we can apply SERENEt is mainly related to limitations in the efficiency of the preprocessing as well as the spin-saturated approximation, $T_{\mathrm{S}}\gg T_{\mathrm{CMB}}$, that we impose on our datasets.
@@
Contrary to our initial intuition, one of our significant discoveries is that training our network on a more realistic prior appears to provide more stable and, in some cases, more accurate predictions. This conclusion was first evident from the results in §4.5, and later it is confirmed in §4.6, where we do a final test and run an additional 21-cm simulation that also provides a galaxy catalogue. From the location of the sources, their redshift and some assumption on the ionizing photons mean-free path, $\lambda_{\mathrm{mfp}}$, we can compute a different binary prior on the assumption that reionization is an inside-out process and thus assuming that a region void of signal, $\delta T_{b}=0\,\rm mK$ is ionized, i.e. $x_{\mathrm{HI}}(\mathbf{r},z)=0$.
@@
There have been several attempts to address foreground residuals in 21-cm observations (e.g. Gagnon-Hartman et al., 2021a; Kennedy et al., 2024). However, these efforts have not focused on reconstructing the complete 21-cm signal image data, which contains valuable non-Gaussian information. Moreover, similar works have combined standard foreground mitigation techniques with neural network approaches. A similar approach has been applied to Post-EoR (Makinen et al., 2021b; Shi et al., 2024). However, only recently Chen et al. (2024) has tested the stability and limitations of these methods to unexpected systematic and different models employed in the training/validation of the network. Our work provides better results than previous works in recovering both Hi and Hii regions at small scales. It also provides a comprehensive statistical analysis of the recovery efficiency across a wide range of astrophysical parameters and redshifts (Li et al., 2019; Sabti et al., 2024), indicating that in some scenarios, neural networks encounter greater difficulties in recovering the 21-cm signal. This latter point requires further investigation and will be part of our future work.
With SERENEt, we can explore higher-order summary statistics, such as the bispectrum (e.g. Majumdar et al., 2018; Watkinson et al., 2019) and improve our understanding of the underlying astrophysical and cosmological processes that affect the signal (e.g. Jennings et al., 2020; Hutter et al., 2020; Giri & Schneider, 2022; Schaeffer et al., 2023; Schneider et al., 2023).